# Ramesh Karri, Ozgur Sinanoglu, Muhammad Shafique, Johann Knechtel NYU Tandon School of Engineering, USA NYU Abu Dhabi, UAE

### v2

@@ Abstract
Large Language Models (LLMs) offer transformative capabilities for hardware design automation, particularly in Verilog code generation. However, they also pose significant data security challenges, including Verilog evaluation data contamination, intellectual property (IP) design leakage, and the risk of malicious Verilog generation. We introduce SALAD, a comprehensive assessment that leverages machine unlearning to mitigate these threats. Our approach enables the selective removal of contaminated benchmarks, sensitive IP and design artifacts, or malicious code patterns from pre-trained LLMs, all without requiring full retraining. Through detailed case studies, we demonstrate how machine unlearning techniques effectively reduce data security risks in LLM-aided hardware design.
@@ II-B Data Security and Privacy of LLMs
For hardware design, these threats are only recently studied to some degreee. RTL-Breaker [38] demonstrates backdoor injection into LLMs to synthesize hardware with malicious triggers. VeriContaminated [3] investigates data contamination in foundational models for Verilog generation. VeriLeaky [4] explores data extraction attacks on fine-tuned LLMs and evaluates logic locking as defense.
@@ III Threat Model and Our Approach
Our workflow is shown in Fig. 2. Sensitive LLMs are models fine-tuned on contaminated, proprietary, malicious, or IP data combined with open-source datasets, reflecting the real-world scenarios. In contrast, clean LLMs are fine-tuned solely on open-source datasets and are assumed free of sensitive information. Using this setup, we investigate three Research Questions (RQs):
@@
See also Appendix -A for more details on the formalism underlying for our approach.
@@ IV Experimental Setup
Sensitive LLMs. We fine-tuned five models with selected datasets: (1) Benchmark contamination from 156 [5] and 50 [6] design challenges; (2) 1,134 custom designs from RTL-Repo [48] test set, collected from public GitHub repositories; (3) 703 secret, in-house IP designs developed through years of applied research and multiple tapeouts [redacted for blind review]; and (4) 835 designs poisoned with payload [redacted for blind review]. Each dataset is combined with the RTL-Coder [1] training dataset. We use LLaMA 3.1-8B as baseline model for all evaluations. Fine-tuning is performed for 3 epochs with a learning rate of 1e-5 using the Adam optimizer. For inference, we set the temperature to 0.8, top-p to 0.75, and maximum context length to 2048 tokens.
Dataset Split: Per standard protocols, dataset splits are:
@@
  • Forget Dataset: As outlined in Section III they include contaminated, custom, proprietary, or malicious data marked for unlearning.
@@
Unlearning Evaluation Metrics. We assess the efficacy of unlearning in LLMs using the following metrics. See Appendix -B for more details on each metrics.
@@
  • Min-Kand Min-K%++: Min-K[50] computes a score by averaging the likelihoods of the klowest-probability tokens. Min-K%++ [51] extends this method calibrating based on token distribution statistics, yielding a robust and theoretically grounded detection approach. We selected Min-K%++ in our experiments.
@@
Method VerilogEval – Pass@K on 156 Samples RTLLM – Syntax/Func Pass@K on on 50 Samples Pass@1 Pass@5 Pass@10 Pass@15 Pass@1 Pass@5 Pass@10 Pass@15 EP1 EP2 EP3 EP1 EP2 EP3 EP1 EP2 EP3 EP1 EP2 EP3 EP1 EP2 EP3 EP1 EP2 EP3 EP1 EP2 EP3 EP1 EP2 EP3 GA 48 3 0 69 18 0 77 24 1 78 27 1 1/0 1/0 1/0 1/0 1/0 1/0 1/0 1/0 1/0 1/0 1/0 1/0 GD 43 29 14 64 38 20 70 40 20 72 40 21 1/0 1/0 4/0 1/0 1/0 11/1 1/0 1/0 18/1 1/0 1/0 24/3 PO 45 37 23 71 63 50 77 69 58 82 72 63 1/0 1/0 1/0 1/0 1/0 1/0 1/0 1/0 1/0 1/0 1/0 1/0 NPO 46 41 33 68 58 50 78 61 55 80 63 57 2/0 2/0 3/0 5/1 8/0 6/1 6/1 10/1 7/2 10/1 10/1 8/2 RMU 31 39 35 56 56 60 61 62 72 63 65 75 21/10 25/9 26/11 29/14 33/15 34/14 34/16 34/15 36/16 34/17 39/16 37/17 SimNPO 35 37 32 59 57 53 68 62 62 72 63 67 20/5 27/12 22/7 32/12 34/16 35/14 36/14 37/17 36/17 37/15 37/17 39/18 Sensitive 43 43 43 65 65 65 74 74 74 82 82 82 24/12 24/12 24/12 35/16 35/16 35/16 35/17 35/17 35/17 37/19 37/19 37/19 Clean 38 38 38 74 74 74 79 79 79 83 83 83 25/9 25/9 25/9 35/14 35/14 35/14 37/16 37/16 37/16 39/18 39/18 39/18
@@ V-B Experiment Results
Unlearning Methods. Fig. 3 compares unlearning algorithms on RTLLM- and VerilogEval-contaminated models. On RTLLM-contaminated models, gradient-based methods (GA, GD) and PO exhibit over-aggressive forgetting, reducing FR from 65.1% to just 1.0–1.9% at EP3. In contrast, RMU and SimNPO demonstrate more controlled unlearning, with FR reduced to 44.4% and 60.3% respectively, while significantly lowering memorization leakage from 53.8% to 22.2% (RMU) and 36.8% (SimNPO). A similar trend is also observed for VerilogEval-contaminated models. Notably, SimNPO achieves the best leakage mitigation, reducing Min-K%++ from 85.1% to 30.8%, while PO demonstrates instability by spiking memorization to 99.9% at EP1.
@@
Unlearning Epochs. We further evaluate unlearning across epochs EP1 to EP3, finding that prolonged training generally intensifies forgetting, but not always desirably. Under RTLLM contamination, PO’s FR drops from 9.7% to 1.9%, but resulting in functional collapse despite persistently high Min-K%++ values. By contrast, RMU exhibits gradual FR reduction (53.1% to 14.4%) alongside consistent Min-K%++ decline (46.1% to 22.2%), reflecting balanced forgetting and stability. For VerilogEval contamination, FR and Min-Kshow unstable behaviour under GD with an anomalous surge in Min-K%++ to 99.9% at EP1, suggesting overfitting to the forget set. SimNPO, however, remains stable across epochs, with FR decreasing (81.3% to 76.5%) and Min-K%++ improving steadily (57.5% to 30.8%), indicating effective unlearning.
@@ VI-A Overview
As the custom design in this study is sourced from GitHub, it faces similar data contamination issues as RTLLM and VerilogEval. To ensure rigorous evaluation, we adopt the setup from [3], using Min-Kand CDD to assess contamination levels under the clean model. Results are shown in Table II. We observe severe data contamination in three open-source benchmarks, with rates up to 100% (Min-K%) and 69.87% (CDD). RTLLM is the most affected, underscoring critical reliability concerns in current evaluation practices.
@@ VI-B Experiment Results
Unlearning Epochs. For cross-epoch analysis over 3 unlearning epoches and the clean model, GD shows immediate complete forgetting by EP1 (39.7% to 0.0%), reflecting aggressive unlearning. SimNPO follows a steadier trajectory (39.7% to 29.0% over 3 epochs), maintaining performance stability. Min-K%++ trends mirror this: GD drives it near-zero by EP1, while RMU and SimNPO converge at 0.7% and 10.2% respectively by EP3. PO shows gradual improvement from 81.8% to 84.2% across epochs. Most methods converge by EP2 or 3, with SimNPO and RMU showing minimal drift after epoch 2. These results suggest EP3 provides an effective balance between forgetting and efficiency.
@@ VIII-B Experiment Results
Unlearning Epochs. Epoch-level trends reveal differing dynamics in in-house IP leakage reduction from the baseline with FR of 45.1% and Min-K%++ of 89.1% over 3 unlearning epochs. Gradient methods forget proprietary content aggressively: GA lowers FR to 0.5% by EP1, stabilizing at 1.7% by EP3; GD reduces FR steadily from 45.1% to 0.5% over 3 epochs with stable Min-K%++ decline from 89.1% to 1.6% to 0.4% with unlearning. GA’s Min-K%++ fluctuates from 96.7% to 53.1% then increases to 59.0%, which shows unsuitable large unlearning epoch setup may cause potential in-house IP recovery issues. Preference-based methods improve gradually: NPO reaches FR 15.0%, Min-K%++ 11.8% by EP3; RMU drops FR early to 8.6% with minor metric noise later. SimNPO initially improves from 45.1% for FR to 0.8%, but Min-K%++ rebounds sharply (from 2.5% to 36.4%). In short, 2–3 epochs suffice for unlearning convergence on in-house IP, possibly due to the high complexity of proprietary IPs used in this evaluation, meaning that higher complexity may not be suitable for longer epochs. GA and GD yield fast, strong in-house IP protection, while preference-based ones require extended training for comparable security.
@@ X Conclusion and discussion
Based on our observation, RMU and SimNPO reduce Min-K%++ from 85.1% to 30.8% on VerilogEval contamination, demonstrating effective forgetting of sensitive hardware knowledge (RQ1). They achieve stable unlearning within 2–3 epochs, while GA/GD offer more aggressive erasure but degrade utility (RQ2). Despite unlearning, models retain strong RTL generation performance—RMU reaches a Pass@15 of 75 (vs. 83 for clean models), indicating minimal trade-offs (RQ3).
Future work includes designing unlearning algorithms tailored to code generation, developing more rigorous evaluation protocols, and exploring the effect of unlearning on alternative architectures such as reasoning-focused LLMs.

### v3

@@ Abstract
Large Language Models (LLMs) offer transformative capabilities for hardware design automation, particularly in Verilog code generation. However, they also pose significant data security challenges, including Verilog evaluation data contamination, intellectual property (IP) design leakage, and the risk of malicious Verilog generation. We introduce SALAD, a comprehensive assessment that leverages machine unlearning to mitigate these threats. Our approach enables the selective removal of contaminated benchmarks, sensitive IP and design artifacts, or malicious code patterns from pre-trained LLMs, all without requiring full retraining. Through detailed case studies, we demonstrate how machine unlearning techniques effectively reduce data security risks in LLM-aided hardware design. Codes are available at https://github.com/DfX-NYUAD/SALAD.
@@
publicationid: pubid: 979-8-3315-3762-3/25/\$31.00 © 2025 IEEE footnotetext: *Authors contributed equally to this research.footnotetext: 979-8-3315-3762-3/25/\$31.00 © 2025 IEEE
@@ II-B Data Security and Privacy of LLMs
For hardware design, these threats are only recently studied to some degree. RTL-Breaker [38] demonstrates backdoor injection into LLMs to synthesize hardware with malicious triggers. VeriContaminated [3] investigates data contamination in foundational models for Verilog generation. VeriLeaky [4] explores data extraction attacks on fine-tuned LLMs and evaluates logic locking as defense.
@@ III Threat Model and Our Approach
Our workflow is shown in Fig. 2. Sensitive LLMs are models fine-tuned on contaminated, proprietary, malicious, or IP data combined with open-source datasets, reflecting real-world scenarios. In contrast, clean LLMs are fine-tuned solely on open-source datasets and are assumed free of sensitive information. Using this setup, we investigate three Research Questions (RQs):
@@
See also Appendix -A for more details on the formalism underlying our approach.
@@ IV Experimental Setup
Sensitive LLMs. We fine-tuned five models with selected datasets: (1) Benchmark contamination from 156 [5] and 50 [6] design challenges; (2) 1,134 custom designs from RTL-Repo [48] test set, collected from public GitHub repositories; (3) 703 secret, in-house IP designs developed through years of applied research and multiple tapeouts [49, 50, 51, 52, 53, 54, 55]; and (4) 835 designs poisoned with payload [38]. Each dataset is combined with the RTL-Coder [1] training dataset. We use LLaMA 3.1-8B as the baseline model. Fine-tuning is performed for 3 epochs with a learning rate of 1e-5 using the Adam optimizer. For inference, we set the temperature to 0.8, top-p to 0.75, and maximum context length to 2048 tokens.
Dataset Split. Per standard protocols, the dataset is split as follow:
@@
  • Forget Dataset: As outlined in Section III, they include contaminated, custom, proprietary, or malicious data marked for unlearning.
@@
Unlearning Evaluation Metrics. We assess the efficacy of unlearning in LLMs using the following metrics. See Appendix -B for more details on each metric.
@@
  • Min-K% and Min-K%++: Min-K% [57] computes a score by averaging the likelihoods of the k% lowest-probability tokens. Min-K%++ [58] extends this method by calibrating based on token distribution statistics, yielding a robust and theoretically grounded detection approach. We select Min-K%++ in our experiments.
@@
Method VerilogEval – Pass@K on 156 Samples RTLLM – Syntax/Func Pass@K on 50 Samples Pass@1 Pass@5 Pass@10 Pass@15 Pass@1 Pass@5 Pass@10 Pass@15 EP1 EP2 EP3 EP1 EP2 EP3 EP1 EP2 EP3 EP1 EP2 EP3 EP1 EP2 EP3 EP1 EP2 EP3 EP1 EP2 EP3 EP1 EP2 EP3 GA 48 3 0 69 18 0 77 24 1 78 27 1 1/0 1/0 1/0 1/0 1/0 1/0 1/0 1/0 1/0 1/0 1/0 1/0 GD 43 29 14 64 38 20 70 40 20 72 40 21 1/0 1/0 4/0 1/0 1/0 11/1 1/0 1/0 18/1 1/0 1/0 24/3 PO 45 37 23 71 63 50 77 69 58 82 72 63 1/0 1/0 1/0 1/0 1/0 1/0 1/0 1/0 1/0 1/0 1/0 1/0 NPO 46 41 33 68 58 50 78 61 55 80 63 57 2/0 2/0 3/0 5/1 8/0 6/1 6/1 10/1 7/2 10/1 10/1 8/2 RMU 31 39 35 56 56 60 61 62 72 63 65 75 21/10 25/9 26/11 29/14 33/15 34/14 34/16 34/15 36/16 34/17 39/16 37/17 SimNPO 35 37 32 59 57 53 68 62 62 72 63 67 20/5 27/12 22/7 32/12 34/16 35/14 36/14 37/17 36/17 37/15 37/17 39/18 Sensitive 43 43 43 65 65 65 74 74 74 82 82 82 24/12 24/12 24/12 35/16 35/16 35/16 35/17 35/17 35/17 37/19 37/19 37/19 Clean 38 38 38 74 74 74 79 79 79 83 83 83 25/9 25/9 25/9 35/14 35/14 35/14 37/16 37/16 37/16 39/18 39/18 39/18
@@ V-B Experiment Results
Unlearning Methods. Fig. 3 compares unlearning algorithms on RTLLM- and VerilogEval-contaminated models. On RTLLM-contaminated models, gradient-based methods (GA, GD) and PO exhibit over-aggressive forgetting, reducing FR from 65.1% to just 1.0–1.9% at EP3. In contrast, RMU and SimNPO demonstrate more controlled unlearning, with FR reduced to 44.4% and 60.3% respectively, while significantly lowering memorization from 53.8% to 22.2% (RMU) and 36.8% (SimNPO). A similar trend is also observed for VerilogEval-contaminated models. Notably, SimNPO achieves the best leakage mitigation, reducing Min-K%++ from 85.1% to 30.8%, while PO demonstrates instability by spiking memorization to 99.9% at EP1.
@@
Unlearning Epochs. We further evaluate unlearning across epochs EP1 to EP3, finding that prolonged training generally intensifies forgetting, but not always desirably. Under RTLLM contamination, PO’s FR drops from 9.7% to 1.9%, but resulting in functional collapse despite persistently high Min-K%++ values. By contrast, RMU exhibits gradual FR reduction (53.1% to 14.4%) alongside consistent Min-K%++ decline (46.1% to 22.2%), reflecting balanced forgetting and stability. For VerilogEval contamination, FR and Min-K% show unstable behavior under GD with an anomalous surge in Min-K%++ to 99.9% at EP1, suggesting overfitting to the forget set. SimNPO, however, remains stable across epochs, with FR decreasing (81.3% to 76.5%) and Min-K%++ improving steadily (57.5% to 30.8%), indicating effective unlearning.
@@ VI-A Overview
As the custom design in this study is sourced from GitHub, it faces similar data contamination issues as RTLLM and VerilogEval. To ensure rigorous evaluation, we adopt the setup from [3], using Min-K% and CDD to assess contamination levels under the clean model. Results are shown in Table II. We observe severe data contamination in three open-source benchmarks, with rates up to 100% (Min-K%) and 69.87% (CDD). RTLLM is the most affected, underscoring critical reliability concerns in current evaluation practices.
@@ VI-B Experiment Results
Unlearning Epochs. For cross-epoch analysis over 3 unlearning epochs and the clean model, GD shows immediate complete forgetting by EP1 (39.7% to 0.0%), reflecting aggressive unlearning. SimNPO follows a steadier trajectory (39.7% to 29.0% over 3 epochs), maintaining performance stability. Min-K%++ trends mirror this: GD drives it near-zero by EP1, while RMU and SimNPO converge at 0.7% and 10.2% respectively by EP3. PO shows gradual improvement from 81.8% to 84.2% across epochs. Most methods converge by EP2 or 3, with SimNPO and RMU showing minimal drift after epoch 2. These results suggest EP3 provides an effective balance between forgetting and efficiency.
@@ VIII-B Experiment Results
Unlearning Epochs. Epoch-level trends reveal differing dynamics in in-house IP leakage reduction from the baseline with FR of 45.1% and Min-K%++ of 89.1% over 3 unlearning epochs. Gradient methods forget proprietary content aggressively: GA lowers FR to 0.5% by EP1, stabilizing at 1.7% by EP3; GD reduces FR steadily from 45.1% to 0.5% over 3 epochs with stable Min-K%++ decline from 89.1% to 1.6% to 0.4% with unlearning. GA’s Min-K%++ fluctuates from 96.7% to 53.1% then increases to 59.0%, which shows unsuitable large unlearning epoch setup may cause potential in-house IP recovery issues. Preference-based methods improve gradually: NPO reaches FR 15.0%, Min-K%++ 11.8% by EP3; RMU drops FR early to 8.6% with minor metric noise later. SimNPO initially improves from 45.1% for FR to 0.8%, but Min-K%++ rebounds sharply (from 2.5% to 36.4%).
In short, 2–3 epochs suffice for unlearning convergence on in-house IP, possibly due to the high complexity of proprietary IPs used in this evaluation, meaning that higher complexity may not be suitable for longer epochs. GA and GD yield fast, strong in-house IP protection, while preference-based ones require extended training for comparable security.
@@ X Conclusion and discussion
Based on our observations, RMU and SimNPO reduce Min-K%++ from 85.1% to 30.8% on VerilogEval contamination, demonstrating effective forgetting of sensitive hardware knowledge (RQ1). They achieve stable unlearning within 2–3 epochs, while GA/GD offer more aggressive erasure but degrade utility (RQ2). Despite unlearning, models retain strong RTL generation performance—RMU reaches a Pass@15 of 75 (vs. 83 for clean models), indicating minimal trade-offs (RQ3).
Future work includes designing unlearning algorithms for code generation, establishing stricter evaluation protocols, and exploring unlearning on reasoning-focused LLMs.